# An Efficient Learning Control Framework with Sim-to-Real for String-Type Artificial Muscle-Driven Robotic Systems

### v3

@@ Abstract
Reinforcement learning for robot control, artificial muscle, sim-to-real, sample efficiency.
@@ I Introduction
Artificial muscles, designed to mimic biological counterparts, generate mechanical forces and movements in response to external stimuli like electrical or chemical signals. Well known for their high power-to-weight ratios and inherent compliance[1], artificial muscles offer a unique combination of strength and flexibility conducive to versatile and adaptive robot designs. Over the past decade, robots driven by artificial muscles have achieved significant advancements, demonstrating impressive performance across diverse domains, including but not limited to biomimetic robots[2, 3], robot manipulators[4, 5], soft robots[6, 7], medical robot[8], origami robots[9], and marine robots[10].
Previous research efforts have primarily centered around structural design and open-loop motion performance evaluation[2, 3, 4, 5, 6, 7, 8, 9, 10]. Despite their potential, achieving accurate control of string-type artificial muscle-driven robots remains challenging due to their highly nonlinear and complex dynamics. Traditional control strategies, such as proportional-integral-derivative (PID) control[11, 12, 13], energy-based nonlinear control[14], and sliding mode control[15, 16] have been investigated to mitigate these challenges. Although these approaches have demonstrated success in specific tasks, a comprehensive, unified control strategy for string-type artificial muscle-driven robots is still lacking. As deep reinforcement learning (DRL) rapidly develops, learning-based controllers[17, 18, 19, 20, 21] have emerged as a promising alternative, offering the potential to overcome the limitations of conventional methods. The standard procedure of applying DRL in the targeted robotic systems typically follows the schematic illustrated in Fig. 1. Initially, identifying the relevant parameters and establishing a simulation environment enable a safe and efficient training process. The DRL algorithms are then trained within this simulation environment to develop an effective control policy. Finally, the learned control policy is transferred to the physical robotic platform to achieve the desired control objectives.
@@
However, several challenges persist in the applications of DRL to string-type artificial muscle-driven robots. One major issue is the sim-to-real gap, which arises from inaccuracies in dynamic models, system identification errors, and variability in muscle string parameters[22, 23]. These factors often result in control strategies that demonstrate strong performance in simulation but struggle to transfer effectively to real-world applications. Another challenge is the low learning efficiency of DRL algorithms, which require substantial amounts of training data and iterations due to the complexity of neural networks and the trial and error learning process[24]. Addressing these two challenges is crucial for advancing control technologies in artificial muscle-driven robots, which are increasingly employed in soft robotics and bio-inspired systems.
This paper presents a DRL-based control strategy designed to enhance the data efficiency and real-world applicability of controllers for artificial muscle-driven robots. The proposed framework is evaluated on two robotic systems actuated by artificial muscle strings: a biomimetic robotic eye driven by super-coiled polymers (SCPs)[25, 26] and a parallel robotic wrist actuated by twisted-coiled actuators (TCAs)[27, 28]. Specifically, we introduce a novel augmentation method for state-based data, which transforms a single transition data tuple into several, and we leverage an easy-to-implement PID controller to generate trajectories for early-stage training of DRL agents. Compared to the baseline algorithm (i.e., Soft Actor Critic [29]), the required training data is reduced by approximately $60\%$ in the targeted robotic systems. Additionally, we leverage a zero-shot sim-to-real transfer technique on artificial muscle dynamics. The experimental results demonstrate a significant improvement in real-world control performance with the averaged error reduced by $36.3\%$ through the application of this method.
The remainder of this paper is organized as follows. Section II provides a review of the relevant literature. Section III introduces the basic concepts of applying reinforcement learning to the robotic systems studied. Section IV offers a concise overview of the robotic platforms utilized in this study. In Section V, we present a detailed description of our proposed efficient learning control framework. Section VI and Section VII discuss the experimental results, including training outcomes in simulation and control performances in real-world scenarios, respectively. Finally, concluding remarks and current limitations are discussed in Section VIII.
@@ II Related Work
The control of artificial muscle-driven robotic systems using learning-based methods has garnered significant attention in recent years[30, 31, 20, 19, 32, 21, 18, 33]. Among these approaches, DRL has demonstrated effectiveness in handling diverse and complex tasks, positioning it as a promising control solution for a wide range of applications. An increasing number of researchers have leveraged DRL algorithms to address control challenges in artificial muscle-driven robots, spanning applications ranging from dexterous grasping of robotic hands[17] to agile motion control of soft robotic manipulators[33, 18], and other related applications[32, 21]. Many of these studies have noted that the low sample efficiency often leads to prolonged training time[20, 21], but few have focused specifically on improving training efficiency. Several existing state-of-the-art (SOTA) reinforcement learning (RL) methods may be applicable[34, 35, 36, 37], but they have not been evaluated within the context of the robotic systems we are addressing. The unique nonlinear and slow response characteristics of artificial muscle may present challenges that are not adequately addressed by these existing methods. Our previous work[19] explored low learning efficiency problem but was limited to simulation-based experiments.
Due to the sim-to-real gap, control strategies that perform well in simulation may not seamlessly transfer to real-world applications[33, 18]. Despite being labor-intensive, some researchers[31, 32] have attempted to directly train DRL algorithms on real robots. However, the artificial muscles considered in this paper are all thermally actuated, which inherently leads to energy inefficiency and limited cycling rates[38]. Continuous rapid transitions between heating and cooling often cause irreversible damage to the muscles, making direct training on real robots unsafe. It is worth noting that in recent years, there have been some efforts to bridge the sim-to-real gap of artificial muscles. Many researchers[22, 23, 39] have proposed more accurate yet significantly more complex dynamic models. In this paper, however, we aim to directly improve the performance of DRL-based controllers by employing a sim-to-real transfer technique, thereby reducing the need for complicated mathematical modeling and precise parameter identification for artificial muscles.
@@ III Preliminaries
  • $\mathcal{R}$ is the reward function that maps the current state $s_{t}$ and action $a_{t}$ to an immediate reward $r_{t}=\mathcal{R}(s_{t},a_{t})$. The design of reward function plays a vital role in the learning process of the DRL agents, which learn the optimal policy under the guidance thereby. In this paper, we develop a generalized quadratic-form reward function designed for the robotic systems of interest:
@@
    where $e_{t}=|\bm{y}^{*}-\bm{y}_{t}|$ is the system tracking error, and $Q_{e}$ and $R_{a}$ represent the output and control weight matrices, respectively. A bonus reward $r_{\text{bonus}}$ is given when the system tracking error falls below a sufficiently small threshold.
@@ IV Experimental Platforms
This section presents an overview of the actuators and robotic platforms used in this paper, including SCP and TCA artificial muscles (Fig. 2), a 2-DOF robotic eye (Fig. 3) and a parallel robotic wrist (Fig. 4). For each system, we briefly describe the hardware design, state space, action space, and reward function.
@@ IV-B Robotic Eye
The high occurrence of unsuccessful corrective surgeries for ocular motor disorders underscores the need for a muscle-driven robotic eye platform to assist ophthalmologists in gaining deeper insights into the biomechanics of human eye movement. In our previous work[25], we proposed a prototype of a 2-DOF robotic eye and derived the geometric and dynamic models specific to this prototype. As illustrated in Fig. 3, the hemispherical eyeball is driven by two pairs of SCP muscle actuators, accommodating a Raspberry Pi Mini Camera (RPiMC). The primary control objective is to precisely and rapidly adjust the eyeball’s orientation towards the desired target angles by regulating the SCP muscles, namely foveation control in the eye movement study.
@@
where $\theta,\dot{\theta},\psi,\dot{\psi},$ represent the pitch, yaw angles, and their respective angular velocities for control. $\Delta T_{i}$ represents the temperature difference between the ambient temperature $T_{0}$ and the current temperature $T_{i}$ of muscle $m_{i}$.
@@
Here, $\theta^{*}$ and $\psi^{*}$ represent the target angles. The Euler angles and corresponding angular velocities are measured through visual feedback from the embedded RPiMC. The actuating voltage applied to muscle $m_{i}$ is denoted as $V_{i}$. Given the antagonistic configuration of the muscle pairs, the control input is defined by a constrained action vector, i.e.,
@@
The constraint requires that only one muscle is activated at each time step in the pairs $(m_{1},m_{2})$ and $(m_{3},m_{4})$. The reward function follows the structure of (1), where the output and control weight matrices were set to $Q_{e}=\text{diag}\{0.05,0.25,0.05,0.25\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively. They were chosen based on empirical tuning to balance the trade-off between output error and control effort. The bonus reward is granted when the tracking error is sufficiently small, given by
@@ IV-C Robotic Wrist
Research into artificial muscle-driven robotic wrist joints is of great importance to several domains, including the development of dexterous robotic systems with human-like capabilities, advancements in rehabilitation and prosthetics technology, and the enhancement of robotic manipulation versatility across various fields. We designed a parallel robotic wrist prototype driven by three equally spaced compound TCAs[27] as illustrated in Fig. 4. The base plate remains stationary, while the upper plate, also known as the end effector, moves via three muscle strings. Linkages between the two plates are constructed to support a two degree-of-freedom motion of the upper plate[40]. Given this configuration, we developed a simplified dynamics model under the assumption of a fixed rotational center for the upper plate. Similar to the foveation control of the robotic eye, the control objective entails precise regulation of the moving upper plate’s orientation. This necessitates adjusting the applied actuating voltages across the three muscle strings to drive the upper plate from its initial equilibrium state to predefined target angles.
@@
where $\theta^{*}$ and $\phi^{*}$ denote the target angles, with the respective Euler angles and angular velocities measured using the inertial measurement unit (IMU) positioned at the center of the upper plate. Additionally, the input action vector $a_{t}$ of the system is defined as:
@@
The output and control weight matrices of the reward function are set to $Q_{e}=\text{diag}\{0.05,0.2,0.05,0.2\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively. When the tracking error is sufficiently small, a bonus reward is granted, i.e.,
@@ V Learning Control Framework
This section provides a comprehensive overview of the proposed learning control framework, as illustrated in Fig. 5. The soft actor-critic (SAC) algorithm serves as the baseline DRL method. To improve data efficiency, we employ a two-stage training strategy. Initially, a standard PID controller is initialized to generate imperfect trajectories for the DRL agent. Subsequently, the agent interacts directly with the simulation environment to update its policy. All transition data collected from the environment undergoes a state-based augmentation strategy before being stored into the replay buffer. To mitigate the sim-to-real gap, we introduce a randomization method on the artificial muscle dynamics at the beginning of each training episode. Eventually, the trained actors (control policies) can be directly implemented in real-world robotic systems to achieve the desired control performance.
@@ V-A Soft Actor-Critic
Soft Actor-Critic (SAC) [29], a state-of-the-art DRL algorithm, is widely adopted as a strong baseline for continuous control tasks in robotics due to its robustness, sample efficiency, and model-free nature. Unlike traditional DRL algorithms, SAC incorporates an entropy-maximizing approach, with its learning objective defined as:
@@
where $\mathcal{H}(\pi(\cdot|s_{t}))$ represents the entropy of the policy $\pi$, and $\alpha$ serves as a temperature parameter. This entropy-regularized objective improves the robustness of DRL and accelerates its training [24]. To process sequential information critical for addressing partial observability issue and implementing dynamics randomization, we integrate a Gated Recurrent Unit (GRU) [41] as the hidden layer in the actor-critic architecture.
@@ V-B Randomization of Muscle Dynamics
$c$ $(\text{N}/^{\circ}\text{C})$ $0.0055$ uniform$(0.85,1.15)$
$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $0.28$ uniform$(0.8,1.2)$
$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.094$ uniform$(0.85,1.15)$
$R$ $(\Omega)$ $20.0$ uniform$(0.9,1.1)$
$\theta,\psi$ noise $(^{\circ})$ normal$(0,0.1)$
$\dot{\theta},\dot{\psi}$ noise $(^{\circ}/\text{s})$ normal$(0,0.05)$
@@
where $x$ and $x_{0}$ denote the loaded length and unloaded length of the actuator, respectively, while $k$, $b$, and $c$ denote the mean stiffness, damping, and temperature effect coefficients of the actuator, respectively. $T(t)$ denotes the current actuator temperature, while $T_{\text{amb}}$ stands for the ambient temperature of the environment. The thermoelectric model of the actuator is expressed as:
@@
where $C_{\text{th}}$ represents the thermal mass of the actuator (measured in $\text{Ws}/^{\circ}\text{C}$), $P(t)$ indicates the heat applied to the actuator, $\lambda$ is the absolute thermal conductivity of the actuator with respect to the ambient environment (in $\text{W}/^{\circ}\text{C}$). The applied power, $P(t)$, is controlled through the voltage applied to the muscle, expressed as $P(t)=V(t)^{2}/R$.
@@
$c$ $(\text{N}/^{\circ}\text{C})$ $0.0707$ uniform$(0.85,1.15)$
$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $3.06$ uniform$(0.8,1.2)$
$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.1189$ uniform$(0.85,1.15)$
$R$ $(\Omega)$ $10.0$ uniform$(0.9,1.1)$
$\theta,\phi$ noise $(^{\circ})$ normal$(0,0.1)$
$\dot{\theta},\dot{\phi}$ noise $(^{\circ}/\text{s})$ normal$(0,0.05)$
The core idea behind dynamics randomization is to introduce randomness into specific dynamic parameters within the simulator during training. This approach allows the DRL agent to learn a policy adaptable to various dynamics conditions, including those encountered in real-world experiments. Specifically, we randomize several critical muscle parameters within the simulated dynamics environment. A distribution $\rho_{\mu}$ of the muscle parameter set $\mu$ is established to encompass the variability of muscle dynamic parameters for dynamics randomization purposes. This approach is expected to be particularly effective in our setup, as artificial muscles are known to exhibit significant variability due to factors such as material degradation, slight manufacturing differences, and environmental temperature changes. According to the muscle models (11) and (12), the dynamic parameter set $\mu$ for randomization is given by
$\mu:=\left\{k,b,c,C_{\text{th}},\lambda,R\right\}.$ (13)
Table I and II outline the distribution $\rho_{\mu}$ of SCP and TCA muscles, respectively, along with the observation noise added for the robotic eye and wrist systems. $\rho_{\mu}$ is centered on the mean values derived from system identification to align the data distribution in simulation more accurately with that in reality. To strike the right balance between randomization and fidelity, we carefully tune the range and distribution of randomized parameters based on empirical observations of muscle behavior in real-world scenarios. Specifically, the range of each parameter is determined based on the distribution observed across different real muscle strings, as outlined in [26]. This approach allows us to maintain a high degree of fidelity while simultaneously enabling the controller to generalize effectively to real-world uncertainties. Details on the impact of varying randomization ranges on experimental results are discussed in Section VII-C.
During the training process, the physical parameters of the actuators are randomized within predefined scaling ranges (i.e., $\mu\sim\rho_{\mu}$) at the start of each training episode and remain constant throughout. Additionally, Gaussian noise is added to the policy’s observation inputs to simulate the noise typically present in real-world sensor data. This observation noise varies at each time step and across different state variables. Although this study employs a simplified dynamic model for artificial muscles, the proposed method can be extended to more complex muscle dynamics models to further enhance control performance.

### v4

@@ Abstract
Reinforcement learning for robot control, artificial muscle, sample efficiency, sim-to-real.
@@ I Introduction
Artificial muscles, designed to mimic biological counterparts, generate mechanical forces and movements in response to external stimuli like electrical or chemical signals. Well known for their high power-to-weight ratios and inherent compliance[1], artificial muscles offer a unique combination of strength and flexibility conducive to versatile and adaptive robot designs. Over the past decade, robots driven by artificial muscles have achieved significant advancements, demonstrating impressive performance across diverse domains, including but not limited to biomimetic robots[2, 3], robot manipulators[4, 5], and soft robots[6, 7].
Previous research efforts have primarily centered around structural design and open-loop motion performance evaluation[2, 3, 4, 5, 6, 7]. Despite their potential, achieving accurate control of string-type artificial muscle-driven robots remains challenging due to their highly nonlinear and complex dynamics. Traditional control strategies, such as proportional-integral-derivative (PID) control[8, 9, 10], energy-based nonlinear control[11], and sliding mode control[12, 13] have been investigated to mitigate these challenges. Although these approaches have demonstrated success in specific tasks, a comprehensive, unified control strategy for string-type artificial muscle-driven robots is still lacking. As deep reinforcement learning (DRL) rapidly develops, learning-based controllers[14, 15, 16, 17, 18] have emerged as a promising alternative, offering the potential to overcome the limitations of conventional methods. The standard procedure of applying DRL in the targeted robotic systems typically follows the schematic illustrated in Fig. 1. Initially, identifying the relevant parameters and establishing a simulation environment enable a safe and efficient training process. The DRL algorithms are then trained within this simulation environment to develop an effective control policy. Finally, the learned policy is transferred to the physical robotic platform to achieve desired control performance.
@@
However, several challenges persist in the applications of DRL to string-type artificial muscle-driven robots. One major issue is the sim-to-real gap, which arises from inaccuracies in dynamic models, system identification errors, and variability in muscle string parameters[19, 20]. These factors often result in control strategies that demonstrate strong performance in simulation but struggle to transfer effectively to real-world applications. Another challenge is the low learning efficiency of DRL algorithms, which require substantial amounts of training data and iterations due to the complexity of neural networks and the trial and error learning process[21]. Addressing these two challenges is crucial for advancing control technologies in artificial muscle-driven robots.
This paper presents a DRL-based control strategy designed to enhance the data efficiency and real-world applicability of controllers for artificial muscle-driven robots. The proposed framework is evaluated on two robotic systems actuated by artificial muscle strings: a biomimetic robotic eye driven by super-coiled polymers (SCPs)[22, 23] and a parallel robotic wrist actuated by twisted-coiled actuators (TCAs)[24, 25]. Specifically, we introduce a novel augmentation method for state-based data, which transforms a single transition data tuple into several, and we leverage an easy-to-implement PID controller to generate trajectories for early-stage training of DRL agents. Compared to the baseline algorithm (i.e., Soft Actor Critic [26]), the required training data is reduced by approximately $60\%$ in the targeted robotic systems. Additionally, we leverage a zero-shot sim-to-real transfer technique on artificial muscle dynamics. The experimental results demonstrate a significant improvement in real-world control performance with the averaged error reduced by $36.3\%$.
The main contributions of this paper are threefold. First, we enhance the training of DRL by leveraging PID controller-generated trajectories to accelerate early-stage training and introducing a novel state-based augmentation method that transforms single data transitions into multiple augmented samples. Second, we propose a zero-shot sim-to-real transfer technique specifically tailored for artificial muscle dynamics, effectively addressing the sim-to-real gap caused by dynamic model inaccuracies and parameter variability. Third, we evaluate our approach through extensive experiments on two distinct robotic systems, demonstrating significant reductions in both training data requirements and real-world control errors. These results validate the effectiveness of the proposed method and highlight its potential applicability to similar systems.
@@ II Related Work
The control of artificial muscle-driven robotic systems using learning-based methods has garnered significant attention in recent years[27, 28, 17, 16, 29, 18, 15, 30]. Among these approaches, DRL has demonstrated effectiveness in handling diverse and complex tasks, positioning it as a promising control solution for a wide range of applications. An increasing number of researchers have leveraged DRL algorithms to address control challenges in artificial muscle-driven robots, spanning applications ranging from dexterous grasping of robotic hands[14] to agile motion control of soft robotic manipulators[30, 15], and other related applications[29, 18]. Many of these studies have noted that the low sample efficiency often leads to prolonged training time[17, 18], but few have focused specifically on improving training efficiency. Several existing state-of-the-art (SOTA) data efficient reinforcement learning (RL) methods may be applicable[31, 32, 33, 34], but they have not been evaluated within the context of the robotic systems we are addressing. The unique nonlinear and slow response characteristics of artificial muscle may present challenges that are not adequately addressed by these existing methods. Our previous work[16] explored low learning efficiency problem but was limited to simulation-based experiments.
Due to the sim-to-real gap, control strategies that perform well in simulation may not seamlessly transfer to real-world applications[30, 15]. Despite being labor-intensive, some researchers[28, 29] have attempted to directly train DRL algorithms on real robots. However, the artificial muscles considered in this paper are all thermally actuated, which inherently leads to energy inefficiency and limited cycling rates[35]. Continuous rapid transitions between heating and cooling often cause irreversible damage to the muscles, making direct training on real robots unsafe. Many researchers[19, 20, 36] have proposed more accurate yet significantly more complex dynamic models to bridge the sim-to-real gap of artificial muscles. In this paper, however, we aim to directly improve the performance of DRL-based controllers by employing a sim-to-real transfer technique, thereby reducing the need for complicated mathematical modeling and precise parameter identification for artificial muscles.
@@ III Preliminaries
  • $\mathcal{R}$ is the reward function that maps the current state $s_{t}$ and action $a_{t}$ to an immediate reward $r_{t}=\mathcal{R}(s_{t},a_{t})$. In this paper, we develop a generalized quadratic-form reward function designed for the robotic systems of interest:
@@
    where $e_{t}=|\bm{y}^{*}-\bm{y}|$ is the system tracking error, and $Q_{e}$ and $R_{a}$ represent the output and control weight matrices, respectively. These weights are determined based on empirical considerations with diagonal elements set to ensure terms in the reward function have comparable magnitudes. A bonus reward $r_{\text{bonus}}$ is given when the tracking error falls below a sufficiently small threshold.
@@ IV Experimental Platforms
This section presents an overview of the actuators and robotic platforms used in this paper, including SCP and TCA artificial muscles (Fig. 2), a two-degrees-of-freedom (DOF) robotic eye (Fig. 3) and a parallel robotic wrist (Fig. 4). For each system, we briefly describe the hardware design, state space, action space, and reward function design.
@@ IV-B Robotic Eye
The high occurrence of unsuccessful corrective surgeries for ocular motor disorders underscores the need for a muscle-driven robotic eye platform to assist ophthalmologists in gaining deeper insights into the biomechanics of human eye movement. In our previous work[22], we proposed a prototype of a 2-DOF robotic eye and derived the geometric and dynamic models specific to this prototype. As illustrated in Fig. 3, the hemispherical eyeball is driven by two pairs of SCP muscle actuators, accommodating a Raspberry Pi Mini Camera (RPiMC). Here, $Q_{i}$ represents the anchor position of muscle $m_{i}$ on the eyeball, through which the force generated by SCP muscle is transmitted. $P_{i}$ denotes the fixed endpoint of muscle $m_{i}$. The primary control objective is to precisely and rapidly adjust the eyeball’s orientation towards the desired target angles, namely foveation control in eye movement study.
@@
where $\theta,\dot{\theta},\psi,\dot{\psi},$ represent the pitch, yaw angles, and their respective angular velocities for control. $\Delta T_{i}$ represents the temperature difference between the ambient temperature $T_{\text{amb}}$ and the current temperature $T_{i}$ of muscle $m_{i}$.
@@
Here, $\theta^{*}$ and $\psi^{*}$ represent the target angles. The Euler angles and corresponding angular velocities are measured through visual feedback from the embedded RPiMC. In this paper, the actuating voltage applied to muscle $m_{i}$, denoted as $V_{i}$, is treated as a continuous variable within the defined range. Given the antagonistic configuration of the muscle pairs, the control input is defined by a constrained action vector, i.e.,
@@
The constraint requires that only one muscle is activated at each time step in the pairs $(m_{1},m_{2})$ and $(m_{3},m_{4})$. The reward function follows the structure of (1), where the output and control weight matrices are set to $Q_{e}=\text{diag}\{0.05,0.25,0.05,0.25\}$ and $R_{a}=\text{diag}\{0.01,0.01\}$, respectively. The bonus reward is given by
@@ IV-C Robotic Wrist
Research into artificial muscle-driven robotic wrist joints is of great importance to several domains, including the development of dexterous robotic systems with human-like capabilities, advancements in rehabilitation and prosthetics technology, and the enhancement of robotic manipulation versatility across various fields. We design a parallel robotic wrist prototype driven by three equally spaced compound TCAs[24] as illustrated in Fig. 4. The base plate remains stationary, while the upper plate, also known as the end effector, moves via three muscle strings. Linkages between the two plates are constructed to support a two degree-of-freedom motion of the upper plate[37]. Given this configuration, we develop a simplified dynamics model under the assumption of a fixed rotational center for the upper plate. Similar to the foveation control of the robotic eye, the control objective entails precise regulation of the moving upper plate’s orientation.
@@
where $\theta^{*}$ and $\phi^{*}$ denote the target angles, with the respective Euler angles and angular velocities measured using the inertial measurement unit (IMU) positioned at the center of the upper plate. The input action vector $a_{t}$ of the system is defined as:
@@
The weight matrices are set to $Q_{e}=\text{diag}\{0.05,0.2,0.05,0.2\}$ and $R_{a}=\text{diag}\{0.01,0.01,0.01\}$, respectively. The bonus reward is given by
@@
The threshold values in (5) and (9) are selected empirically based on testing experiments. Lower values may result in slower convergence and lower training stability due to reduced positive reinforcement, while higher values could diminish the precision of the learned control policy.
@@ V Learning Control Framework
This section provides a comprehensive overview of the proposed learning control framework, as illustrated in Fig. 5. The Soft Actor-Critic (SAC) algorithm serves as the baseline DRL method. To improve data efficiency, we employ a two-stage training strategy. Initially, a standard PID controller is initialized to generate imperfect trajectories for the DRL agent. Subsequently, the agent interacts directly with the simulation environment to update its policy. All transition data collected from the environment undergoes a state-based augmentation strategy before being stored into replay buffer. To mitigate the sim-to-real gap, we introduce a randomization method on the artificial muscle dynamics. In this paper, we denote the bootstrap method by letter ‘B’, the augmentation method by letter ‘A’, and the randomization of muscle dynamics by letter ‘R’. Thus, our proposed SAC algorithm with all three improvements is denoted by SAC-BAR.
@@ V-A Soft Actor-Critic
SAC [26] is widely adopted as a strong baseline for continuous control tasks in robotics due to its robustness and sample efficiency. SAC incorporates an entropy-maximizing approach, with its learning objective defined as:
@@
where $\mathbb{E}_{(s_{t},a_{t})\sim\rho_{\pi}}$ represents the expectation over the state-action pairs $(s_{t},a_{t})$ sampled from the policy distribution $\rho_{\pi}$, $\mathcal{H}(\pi(\cdot|s_{t}))$ represents the entropy of the policy $\pi$, and $\alpha$ serves as a temperature parameter. This entropy-regularized objective improves the robustness of DRL and accelerates its training [21]. We integrate a Gated Recurrent Unit (GRU) [38] as the hidden layer in the actor-critic architecture to process sequential information.
@@ V-B Randomization of Muscle Dynamics
$c$ $(\text{N}/^{\circ}\text{C})$ $0.0055$ $\mathrm{uniform}(0.85,1.15)$
$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $0.28$ $\mathrm{uniform}(0.8,1.2)$
$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.094$ $\mathrm{uniform}(0.85,1.15)$
$R$ $(\Omega)$ $20.0$ $\mathrm{uniform}(0.9,1.1)$
$\theta,\psi$ noise $(^{\circ})$ $\mathrm{normal}(0,0.1)$
$\dot{\theta},\dot{\psi}$ noise $(^{\circ}/\text{s})$ $\mathrm{normal}(0,0.05)$
@@
where $x$ and $x_{0}$ denote the current length and resting length of the actuator, respectively, while $k$, $b$, and $c$ denote the mean stiffness, damping, and temperature effect coefficients of the actuator, respectively. $T(t)$ denotes the current actuator temperature, while $T_{\text{amb}}$ stands for the ambient temperature of the environment. The thermoelectric model of the actuator is expressed as:
@@
where $C_{\text{th}}$ represents the thermal mass of the actuator (measured in $\text{Ws}/^{\circ}\text{C}$), $P(t)$ indicates the heat applied to the actuator, $\lambda$ is the absolute thermal conductivity of the actuator with respect to the ambient environment (in $\text{W}/^{\circ}\text{C}$). The applied power, $P(t)$, is controlled through the voltage applied to the muscle, expressed as $P(t)=V(t)^{2}/R$, where $R$ is the resistance of the artificial muscle. Equations (11) and (12) describe the dynamics of the artificial muscles. In the simulation, when the RL agent generates an action $a_{t}$, representing the voltages applied to the muscles, these equations are numerically solved to update the muscle states and compute their output forces. The computed forces are then incorporated into the dynamic model of the robotic system to update the overall system state ($s_{t}$).
@@
$c$ $(\text{N}/^{\circ}\text{C})$ $0.0707$ $\mathrm{uniform}(0.85,1.15)$
$C_{\text{th}}$ $(\text{Ws}/^{\circ}\text{C})$ $3.06$ $\mathrm{uniform}(0.8,1.2)$
$\lambda$ $(\text{W}/^{\circ}\text{C})$ $0.1189$ $\mathrm{uniform}(0.85,1.15)$
$R$ $(\Omega)$ $10.0$ $\mathrm{uniform}(0.9,1.1)$
$\theta,\phi$ noise $(^{\circ})$ $\mathrm{normal}(0,0.1)$
$\dot{\theta},\dot{\phi}$ noise $(^{\circ}/\text{s})$ $\mathrm{normal}(0,0.05)$
The core idea behind dynamics randomization is to introduce randomness into specific dynamic parameters within the simulator during training. While this method is typically used in the field of manipulation and locomotion[39, 40], its integration into the control of artificial muscle-driven robots is largely underexplored. We hypothesize that this method is particularly effective due to the unique uncertainties inherent in these systems. Artificial muscle actuators are subject to significant parameter variability caused by factors such as manufacturing tolerances, wear, and environmental conditions. By exposing the RL agent to a diverse range of dynamics during training, this approach facilitates the development of robust control policies that generalize effectively to real-world conditions with unmodeled dynamics.
Specifically, a distribution $\rho_{\mu}$ of the muscle parameter set $\mu$ is established to encompass the variability of muscle dynamic parameters for dynamics randomization purposes. According to (11) and (12), the dynamic parameter set $\mu$ for randomization is given by $\mu:=\left\{k,b,c,C_{\text{th}},\lambda,R\right\}$, where :=assign:=:= is used to denote an assignment operation in this paper. Table I and II outline the distribution $\rho_{\mu}$ of SCP and TCA muscles, respectively, along with the observation noise added for the robotic eye and wrist systems, where $\mathrm{uniform}(a,b)$ denotes a uniform distribution over interval $[a,b]$ and $\mathrm{normal}(a,b)$ denotes a normal distribution with mean $a$ and standard deviation $b$. The distribution $\rho_{\mu}$ is centered on the mean values derived from system identification to ensure alignment with the physical systems. The randomization ranges of these parameters are determined empirically to capture realistic variability while ensuring physical feasibility. These ranges are derived from standard deviations reported in prior studies on artificial muscle properties, including influences from material properties, environmental factors, and manufacturing differences[23, 25, 19]. Details on the impact of randomization variance on control performance are discussed in Section VII-C. During the training process, the physical parameters of the actuators are randomized within predefined scaling ranges (i.e., $\mu\sim\rho_{\mu}$) at the start of each episode and remain constant throughout. Observation noise is sampled independently at each time step from a Gaussian distribution with a constant mean and variance, as specified in Table I and II.